find all zeros of the following polynomial write the polynomial in factored form
f(x) = 7x^3-x^2+7x-1

Answer by lwsshak3(11628)

find all zeros of the following polynomial write the polynomial in factored form
f(x) = 7x^3-x^2+7x-1
factor by grouping
f(x) =x^2(7x-1)+(7x-1)
f(x) = (7x-1)(x^2+1)=0
7x-1=0
x=1/7
or
x^2+1=0
x^2=-1
x=±√-1=±
zeros: 1/7 and two non-real zeros, ±i